# Latent Speech-Text Transformer

### v1

@@ Abstract
Auto-regressive speech-text models are typically pre-trained on a large number of interleaved sequences of text tokens and raw speech encoded as speech tokens using vector quantization. These models have demonstrated state-of-the-art performance in speech-to-speech understanding and generation benchmarks, together with promising scaling laws, primarily enabled by the representational alignment between text and speech. Nevertheless, they suffer from shortcomings, partly owing to the disproportionately longer sequences of speech tokens in contrast to textual tokens. This results in a large compute imbalance between modalities during pre-training as well as during inference, and a potential hindrance to effectively aligning speech and text, ultimately translating to several orders of magnitude slower scaling laws. We introduce the Latent Speech-Text Transformer (LST), which makes pre-training speech-text models more data-efficient by dynamically and inexpensively aggregating speech tokens into latent speech patches. These patches serve as higher-level units that can either align with corresponding textual units to aid capability transfer or even encapsulate common speech sequences like silences to be more compute-efficient. We show that LST outperforms vanilla approaches on speech-to-speech as well as text-to-text benchmarks in both data- and compute-controlled settings, the former indicating more effective representational alignment and the latter indicating steeper scaling laws for speech-text models. On HellaSwag story completion, LST achieves 6.5% absolute gain in speech accuracy under compute-controlled training and 5.3% under data-controlled training, while also improving text performance. We will release our models, code, and the evaluation data to facilitate further research.
@@ 1 Introduction
Inspired by the strong zero- and few-shot understanding and generation capabilities of large auto-regressive textual language models with billions of parameters that are pre-trained on trillions of tokens, Lakhotia et al. (2021) introduce the task of Generative Spoken Language Modeling (GSLM) a.k.a Textless NLP, where raw speech is encoded as a sequence of discrete tokens based on a dictionary of quantized speech features, and an auto-regressive language model (LM) is trained on these tokens with Next Token Prediction (NTP). While initially successful, Cuervo & Marxer (2024) estimate that this approach would require up to three orders of magnitude more data to obtain equivalent capabilities as textual LLMs, largely owing to the same information requiring a significantly larger number of speech tokens to represent compared to text. This increased sequence length also means that these models utilize considerably more compute during inference to process the same amount of semantic content compared to text.
To improve scaling properties of large speech models by taking advantage of the comparatively larger corpus of web text compared to speech, recent efforts have leveraged transfer learning from textual modalities in the form of warm initialization from large pre-trained text models (Hassid et al., 2023), pre-training with interleaved speech-text data (Nguyen et al., 2025), and modeling speech and text in multiple streams to leverage the textual chain of thought or “inner monologue” (Défossez et al., 2024). All these works attempted to some extent to achieve representational alignment between text and speech, where a perfect alignment means the model can treat the two modalities interchangeably without any performance difference. Despite this, there remains a large gap between text-to-text and speech-to-speech performance on the same benchmarks, highlighting the incompleteness of the alignment. We hypothesize that the severe mismatch in information densities between the speech and text tokens is one of the primary factors hindering speech-text alignment.
@@
We first demonstrate that LST models with fixed-size speech patching schemes similar to what Yu et al. (2023) did with text, are able to significantly outperform their non-patching counterparts. Such models are aware of the internals of patches without expending much compute in the process, in contrast with methods that expand the speech token vocabulary by applying subword tokenization, which yield poor downstream performance (Cuervo & Marxer, 2024). We further improve the performance by introducing speech-patching based on textual alignment at the word/subword levels, which crucially also includes patching large sequences of silences. Since this approach requires text-speech alignment timestamps during training and inference, we also introduce a curriculum-based method to eliminate the need for such alignments during inference.
@@
(1) We show improved performance of LST models in both data- and compute-controlled settings compared to vanilla interleaved speech-text models like SpiritLM (Nguyen et al., 2025), as well as models that use subword tokenization, on speech versions of popular text understanding benchmarks such as HellaSwag (Zellers et al., 2019). LST-based models save considerable training and inference compute and improve speech-text representational alignment (see Figure 1).
(2) We introduce different variations of speech patching schemes, including fixed-size static patching and alignment-based patching, and analyze their effectiveness.
(3) We demonstrate that LST continues outperforming the baseline when scaling up the model size from 1B to 7B parameters, which highlights the scalability of our method.
@@ 2 Background
Speech tokenization Approaches for speech tokenization include semantic tokens represented by cluster-ids obtained by k-means clustering of frame representations as in Hubert (Hsu et al., 2021), acoustic tokens obtained as discretized embeddings from residual-vector quantization bottlenecks from self-supervised neural codec models (Zeghidour et al., 2021; Défossez et al., 2024), as well as additional tokens for expressivity and also, combinations of different token categories. In this paper, we follow Hassid et al. (2023); Nguyen et al. (2025) and use Hubert tokens using a codebook of 501 speech tokens at 25Hz. Unlike Nguyen et al. (2025) we do not need to deduplicate Hubert tokens as this is organically handled by the LST architecture.
Sequence Modeling Similar to LLMs for text, speech token modeling is typically done using a large transformer decoder model using causal self-attention, to maximize the likelihood of sequences from a large speech pre-training corpus () in an auto-regressive fashion:
Interleaved Data Since speech sequences are longer and less compact that their corresponding text sequences, such models can require several orders of magnitude more data in order to achieve performance comparable to text models (Cuervo & Marxer, 2024). In order to bridge the gap, Nguyen et al. (2025) find that training on interleaved sequences of text and speech data directly correlates with improved performance. For a subset of the pre-training dataset that contains the textual sequence , where text tokens are obtained using a tokenizer (we use the Llama 2 tokenizer (Touvron et al., 2023) in this paper) and each text token can correspond to a span of speech tokens, the model is trained on an interleaved sequence obtained by replacing arbitrary spans of speech tokens in the sequence sequence with text tokens separated by special modality tokens. This allows the same model to be used for SS, ST, TS and TT tasks. We discuss the process of producing interleaved data from parallel text-speech data in Appendix A.1.
@@ 3 Latent Speech-Text Transformers
The core idea of the LST architecture is to auto-regressively model latent patches of tokens (using a global transformer), rather than individual tokens, similar in spirit to BLT (Pagnoni et al., 2024) which models dynamic-sized patches of bytes. The transformation of speech/text spans to patches and vice-versa takes place with the help of a light-weight local encoder and local decoder, and the entire model is trained end-to-end using the same token-level likelihood as before. Figure 2 illustrates this architecture specialized to the task of speech-text modeling. The majority of the compute expended in terms of FLOPs is in the global transformer, which yields savings by operating on information-dense speech patches instead of granular speech tokens.
Local Encoder. Similar to BLT, the local encoder uses a series of sliding window self-attention and cross-attention layers to aggregate token representations into patch representations. In LST, we only patch spans of speech tokens using strategies described in Section 3.1. Note that a simple alternative to patching is to use subword tokenization methods like Byte Pair Encoding (BPE) on the speech tokens. This was also explored by Cuervo & Marxer (2024) and similar to them, failed to improve performance in our experiments (ablations in Section 5). Unlike BLT, we do not use hash embeddings, as they did not provide improvements in our experiments.
Local Decoder. A light-weight transformer is used as a decoder and trained with NTP loss, with cross-attention layers inserted between every transformer layer. Each token attends to both the previously generated speech patches and text tokens to incorporate patch-level information (using cross-attention) as well as a sliding window of the past 512 tokens (using self-attention).
@@ 3.1 Patching
Let be speech token embeddings obtained using a learned embedding matrix applied to speech tokens . The process of patching maps to a shorter sequence of patch embeddings by aggregating local frame segments. For a frame-index set , a patch embedding is formed via the local encoder:
@@
Static Patching. Speech sequence is split into non-overlapping segments of a fixed length (patch size). Each patch token is obtained by the local encoder from the embeddings in the patch:
For and input embeddings , the first patch is , the second , and so on. Each segment is encoded into a single patch embedding by the local encoder. This provides a uniform compression ratio independent of alignment information.
Alignment Patching. To better synchronize speech and text at the semantic level, alignment patching leverages forced alignment timestamps between speech frames and textual units (e.g. words or BPE tokens). Let denotes the aligned frame ranges, where spans the -th textual unit. The corresponding patch is
Frames outside text spans (e.g., silence) are grouped into separate patches (Fig. 3(a)). For instance, if one word aligns to and the next to , patches are and , with silence forming and .
We obtain alignments with Wav2Vec2+CTC (Baevski et al., 2020), yielding one patch per text unit and silence segment (Fig. 3(b)). While this enforces cross-modal correspondence, it requires an auxiliary model at inference, introducing possible errors. Curriculum patching (Sec. 4) mitigates this by gradually shifting from aligned to static patching during training.
Curriculum Patching. Curriculum patching interpolates between alignment-based and static patching during training. Let denote the probability of using alignment at training step :
@@ 4 Experimental Setup
We next describe the datasets, models, and evaluation protocols used in our experiments.
@@ 4.1 Training and Evaluation Datasets
Text. Our text training data consists of extensive web and academic corpora, sourced from a selected portion of the Llama 2 pre-training collection (Touvron et al., 2023), totaling 1.8T tokens. We follow the LLaMA 2 setup and apply its SentencePiece (Kudo & Richardson, 2018) BPE tokenizer with a 32K vocabulary.
Speech. Our speech training data includes speech which is discretized into HuBERT tokens (501-entry codebook at 25Hz) together with paired text transcriptions. We use LibriLight (60k hours), People’s Speech (30k hours), Multilingual LibriSpeech (50k hours), and Spotify (60k hours), detailed in Table 1. All corpora are aligned using the Wav2Vec2 + CTC framework to provide token-level correspondence between speech and text (Figure 3(b)).
@@
| Dataset | Format | Focus |
| HellaSwag (Zellers et al., 2019) | 1-in-4 MC | Commonsense reasoning |
| StoryCloze (Mostafazadeh et al., 2016) | 1-in-2 MC | Narrative coherence |
| TopicStoryCloze (Hassid et al., 2023) | 1-in-2 MC | Topic consistency |
We omit sWUGGY and sBLiMP (Nguyen et al., 2020), as they target lexical and syntactic judgments on very short speech segments. Such settings are less aligned with our focus on narrative reasoning, where story-level coherence and commonsense understanding are required.
@@ 4.2 LST Models and Baselines
BPE SpeechLLM. Maps speech tokens into 1k BPE units using a SentencePiece tokenizer (Kudo & Richardson, 2018) trained on 100k random speech sequences, replacing speech tokens with BPE-derived units111We use the 1k configuration as our BPE baseline, as larger vocabularies (5k, 10k) showed no benefit..
@@ 5 Results
Compute-controlled. We fix the number of training iterations and per-step sequence budget so that all methods process the same number of units (baseline tokens LST patches). Table 3 shows three trends on HellaSwag. First, patching increases the effective token budget, benefiting both modalities: Curriculum Patching improves TT by +5.2 (47.0→52.2) and SS by +6.5 (39.0→45.5). Second, Aligned Patching is less effective at evaluation, since variable word spans often yield longer patches, reducing the test-time compute. Finally, Mixed and Curriculum patching combine the advantages of shorter evaluation patches with alignment information, consistently outperforming Static and Aligned across datasets.
Data-controlled. Here we fix the data budget with the same amounts of speech and text tokens. Since LST compresses sequences into patches, it processes fewer patch tokens than the baselines, leading to higher efficiency. Table 4 shows that the BPE baseline fails to surpass vanilla SpeechLLM, whereas LST continues to achieve consistent gains. On HellaSwag, Curriculum Patching improves TT accuracy from 49.6 to 52.2 despite reduced computation, while boosting SS from 40.2 to 45.5. Similar improvements are observed on StoryCloze and TopicStoryCloze. Overall, LST with Curriculum Patching reduces the speech–text performance gap from 9.4 to 6.7, demonstrating that alignment through patching benefits both modalities while offering meaningful compute savings.

### v2

@@ Abstract
Auto-regressive speech–text models pre-trained on interleaved text tokens and discretized speech tokens demonstrate strong speech understanding and generation, yet remain substantially less compute-efficient than text LLMs, partly due to the much longer sequences of speech tokens relative to text. This modality imbalance disproportionately allocates pre-training and inference compute to speech, potentially hindering effective cross-modal alignment and slowing performance scaling by orders of magnitude. We introduce the Latent Speech-Text Transformer (LST), which aggregates speech tokens into latent speech patches that serve as higher-level autoregressive units. This design aligns the sequence-modeling granularity between speech and text while improving computational efficiency. The resulting patches can align with textual units to facilitate cross-modal knowledge transfer and compactly capture recurring acoustic patterns such as silence. Across story-completion benchmarks under both compute-controlled and data-controlled settings, LST consistently improves speech accuracy while also improving text performance, achieving up to +6.5% absolute gain on speech HellaSwag in compute-controlled training (+5.3% in data-controlled training). Under compute-controlled scaling from 420M to 1.8B parameters in a near compute-optimal regime, gains grow with scale, and improvements persist up to 7B parameters under fixed-token budgets. These benefits extend to downstream tasks: LST stabilizes ASR adaptation and reduces the effective autoregressive sequence length during ASR and TTS inference, lowering computational cost without degrading reconstruction quality. The code is available at https://github.com/facebookresearch/lst.
@@ 1 Introduction
Inspired by the strong zero- and few-shot understanding and generation capabilities of large auto-regressive textual language models with billions of parameters that are pre-trained on trillions of tokens, Lakhotia et al. (2021) introduce the task of Generative Spoken Language Modeling (GSLM) a.k.a Textless NLP, where raw speech is encoded as a sequence of discrete tokens based on a dictionary of quantized speech features, and an auto-regressive language model (LM) is trained on these tokens with Next Token Prediction (NTP). While initially successful, Cuervo and Marxer (2024) estimate that this approach would require up to three orders of magnitude more data to obtain equivalent capabilities as textual LLMs, largely owing to the same information requiring a significantly larger number of speech tokens to represent compared to text. This increased sequence length also means that these models utilize considerably more compute during inference to process the same amount of semantic content compared to text.
To improve scaling properties of large speech models by taking advantage of the comparatively larger corpus of web text compared to speech, recent efforts have leveraged transfer learning from textual modalities in the form of warm initialization from large pre-trained text models (Hassid et al., 2023), pre-training with interleaved speech-text data (Nguyen et al., 2025), and modeling speech and text in multiple streams to leverage the textual chain of thought or “inner monologue” (Défossez et al., 2024). All these works attempted to some extent to achieve representational alignment between text and speech, where a perfect alignment means the model can treat the two modalities interchangeably without any performance difference. Despite this, there remains a large gap between text-to-text and speech-to-speech performance on the same benchmarks, highlighting the incompleteness of the alignment. We hypothesize that the severe information-density mismatch between speech and text tokens is one of the primary factors hindering speech-text alignment.
@@
We first demonstrate that LST models with fixed-size speech patching schemes similar to what Yu et al. (2023) did with text, are able to significantly outperform their non-patching counterparts. Such models are aware of the internals of patches without expending much compute in the process, in contrast with methods that expand the speech token vocabulary by applying subword tokenization, which yield poor downstream performance (Cuervo and Marxer, 2024). We further improve the performance by introducing speech-patching based on textual alignment at the word/subword levels, which crucially also includes patching large sequences of silences. Since this approach requires text-speech alignment timestamps during training and inference, we also introduce a curriculum-based method to eliminate the need for such alignments during inference.
@@
(1) We show that LST improves performance in both data- and compute-controlled settings compared to prior interleaved speech-text models on speech versions of popular text understanding benchmarks such as HellaSwag (Zellers et al., 2019) (see Figure 1), while substantially reducing training and inference compute and enabling efficient downstream ASR and TTS transfer.
(2) We introduce latent speech patching as a unified mechanism for compressing autoregressive speech sequences and analyze static, alignment-based, and curriculum patching strategies.
(3) We demonstrate that the benefits of LST persist and grow with scale from 1B to 7B parameters, indicating improved sample efficiency and more favorable compute-optimal scaling behavior for spoken language modeling.
Collectively, these contributions demonstrate that aligning the autoregressive modeling granularity of speech and text helps mitigate a key barrier to efficient and scalable spoken language modeling.
@@ 2 Background
Speech tokenization.
Approaches for speech tokenization include semantic tokens represented by cluster-ids obtained by k-means clustering of frame representations as in Hubert (Hsu et al., 2021), acoustic tokens obtained as discretized embeddings from residual-vector quantization bottlenecks from self-supervised neural codec models (Zeghidour et al., 2021; Défossez et al., 2024), as well as additional tokens for expressivity and also, combinations of different token categories. In this paper, we follow Hassid et al. (2023); Nguyen et al. (2025) and use Hubert tokens using a codebook of 501 speech tokens at 25Hz. Unlike Nguyen et al. (2025) we do not need to deduplicate Hubert tokens as this is organically handled by the LST architecture.
Sequence Modeling.
Similar to LLMs for text, speech token modeling is typically done using a large transformer decoder model using causal self-attention, to maximize the likelihood of sequences from a large speech pre-training corpus () in an auto-regressive fashion:
Interleaved Data.
Since speech sequences are longer and less compact that their corresponding text sequences, such models can require several orders of magnitude more data in order to achieve performance comparable to text models (Cuervo and Marxer, 2024). In order to bridge the gap, Nguyen et al. (2025) find that training on interleaved sequences of text and speech data directly correlates with improved performance. For a subset of the pre-training dataset that contains the textual sequence , where text tokens are obtained using a tokenizer (we use the Llama 2 tokenizer (Touvron et al., 2023) in this paper) and each text token can correspond to a span of speech tokens, the model is trained on an interleaved sequence obtained by replacing arbitrary spans of speech tokens in the sequence sequence with text tokens separated by special modality tokens. This allows the same model to be used for SS, ST, TS and TT tasks. We discuss the process of producing interleaved data from parallel text-speech data in Appendix A.1.1.
@@ 3 Latent Speech-Text Transformers
The core idea of the LST architecture is to auto-regressively model latent patches of tokens (using a global transformer), rather than individual tokens, similar in spirit to BLT (Pagnoni et al., 2024) which models dynamic-sized patches of bytes. The transformation of speech/text spans to patches and vice-versa takes place with the help of a light-weight patch encoder and patch decoder, and the entire model is trained end-to-end using the same token-level likelihood as before. Figure 2 illustrates this architecture specialized to the task of speech-text modeling. The majority of the compute expended in terms of FLOPs is in the global transformer, which yields savings by operating on information-dense speech patches instead of granular speech tokens. Latent patching performs bounded-context temporal aggregation, preserving phonetic discriminability while reducing sequence-level redundancy.
Patch Encoder.
Similar to BLT, the patch encoder uses a series of sliding window self-attention and cross-attention layers to aggregate token representations into patch representations. In LST, we only patch spans of speech tokens using strategies described in Section 3.1. Note that a simple alternative to patching is to use subword tokenization methods like Byte Pair Encoding (BPE) on the speech tokens. This was also explored by Cuervo and Marxer (2024) and similar to them, failed to improve performance in our experiments (ablations in Section 5). Unlike BLT, we do not use hash embeddings, as they did not provide improvements in our experiments.
Patch Decoder.
A light-weight transformer is used as a decoder and trained with NTP loss, with cross-attention layers inserted between every transformer layer. Each token attends to both the previously generated speech patches and text tokens to incorporate patch-level information (using cross-attention) as well as a sliding window of the past 512 tokens (using self-attention). Further architectural details of the local patch encoder and decoder, including attention roles and initialization, are provided in Appendix A.2.1.
@@ 3.1 Patching
Let be speech token embeddings obtained using a learned embedding matrix applied to speech tokens . The process of patching maps to a shorter sequence of patch embeddings by aggregating local frame segments. For a frame-index set , a patch embedding is formed via the patch encoder:
@@
Static Patching.
Speech sequence is split into non-overlapping segments of a fixed length (patch size). Each patch token is obtained by the patch encoder from the embeddings in the patch:
For and input embeddings , the first patch is , the second , and so on. Each segment is encoded into a single patch embedding by the patch encoder. This provides a uniform compression ratio independent of alignment information.
Alignment Patching.
To better synchronize speech and text at the semantic level, alignment patching leverages forced alignment timestamps between speech frames and textual units (e.g. words or BPE tokens). Let denotes the aligned frame ranges, where spans the -th textual unit. The corresponding patch is
Frames outside text spans (e.g., silence) are grouped into separate patches (Fig. 3(a)). If consecutive words align to and , patches are and , with silence forming and . We obtain alignments with Wav2Vec2+CTC (Baevski et al., 2020), yielding one patch per text unit and silence segment (Fig. 3(b)). While this enforces cross-modal correspondence, it requires an auxiliary model at inference, introducing possible errors. Curriculum patching (Sec. 3.1) mitigates this by gradually shifting from aligned to static patching during training.
Curriculum Patching.
Curriculum patching gradually transitions from alignment-based to static patching. Let denote the probability of using alignment at training step :
@@ 4.1 Training and Evaluation Datasets
Text.
Speech.
Our speech training data includes speech which is discretized into HuBERT tokens (501-entry codebook at 25Hz) together with paired text transcriptions. We use LibriLight (60k hours), People’s Speech (30k hours), Multilingual LibriSpeech (50k hours), and Spotify (60k hours), detailed in Table 1. All corpora are aligned using the Wav2Vec2 + CTC framework to provide token-level correspondence between speech and text (Figure 3(b)).
@@ 4.2 LST Models and Baselines
BPE SpeechLLM. Maps speech tokens into 1k BPE units using a SentencePiece tokenizer (Kudo and Richardson, 2018) trained on 100k random speech sequences, replacing speech tokens with BPE-derived units111We use the 1k configuration as our BPE baseline, as larger vocabularies (5k, 10k) showed no benefit..
@@ 5 Results
5.1 Performance under Controlled Budgets
Compute-controlled.
We fix the number of training iterations and per-step sequence budget so that all methods process the same number of units (baseline tokens LST patches). Table 3 shows three trends on HellaSwag. First, patching increases the effective token budget, benefiting both modalities: Curriculum Patching improves TT by +5.2 (47.0→52.2) and SS by +6.5 (39.0→45.5). Second, Aligned Patching is less effective at evaluation, since variable word spans often yield longer patches, reducing the test-time compute. Finally, Mixed and Curriculum patching combine the advantages of shorter evaluation patches with alignment information, consistently outperforming Static and Aligned across datasets.
Data-controlled.
Here we fix the data budget with the same amounts of speech and text tokens. Since LST compresses sequences into patches, it processes fewer patch tokens than the baselines, leading to higher efficiency. Table 4 shows that the BPE baseline fails to surpass vanilla SpeechLLM, whereas LST continues to achieve consistent gains. On HellaSwag, Curriculum Patching improves TT accuracy from 49.6 to 52.2 despite reduced computation, while boosting SS from 40.2 to 45.5. Similar improvements are observed on StoryCloze and TopicStoryCloze. Overall, LST with Curriculum Patching reduces the speech–text performance gap from 9.4 to 6.7, demonstrating that alignment through patching benefits both modalities while offering meaningful compute savings.
@@ 5.1 Performance under Controlled Budgets
5.2 Scaling Behavior
Compute-Optimal Scaling.
Figure 4(a) evaluates HellaSwag accuracy under compute-optimal training from 420M to 1.8B parameters. Following Hoffmann et al. (2022), text is trained with 20× model-size tokens, while speech uses half as many tokens to preserve a 2:1 text–speech ratio. At the smallest scale (420M), LST already outperforms the baseline, reaching 29.2% vs. 28.4% for speech and 31.6% vs. 31.1% for text. These improvements compound with scale: at 1.8B, LST (Speech) attains 39.0% compared to 35.3%, while LST (Text) achieves 46.3% over 45.7%. Overall, LST provides consistent gains in both modalities, with advantages apparent from the earliest scale and amplified as model capacity increases.
Sub-Optimal Token Scaling at 7B.
Figure 4(b) shows training dynamics for a 7B model under a fixed processed-token budget (70B tokens), which remains below the scaling-law optimal regime (140B). Under this sub-optimal compute setting, LST exhibits consistently faster improvement and maintains higher accuracy throughout training. Additional comparisons at 1B and 7B are summarized in Appendix A.4, where LST achieves larger gains at 1B and persistent improvements at 7B. Notably, the speech data used in our training has already been reused for multiple epochs (6×), indicating that further token scaling becomes data-limited rather than compute-limited in this regime. The continued upward trend at 7B nevertheless suggests that training toward the scaling-law optimum with additional data would likely further amplify LST’s advantage.
| Task | Model | Iters | clean | other | Ctx. Units | Gen. Units |
| ASR (WER %) | Baseline | 1k | 140 | 202 | 1.0 | – |
| 2k | 44.7 | 73.2 | ||||
| 4k | 20.7 | 42.4 | ||||
| LST | 1k | 6.8 | 10.4 | 0.25 | – | |
| 2k | 6.0 | 13.3 | ||||
| TTS (CER %) | Baseline | 20k | 14.1 | 15.1 | – | 1.0 |
| LST | 20k | 14.1 | 16.2 | 0.25 |
5.3 Downstream Transfer and Analysis
ASR Adaptation and Efficient TTS Generation.
To assess ASR, we fine-tune both models on LibriSpeech clean for 1k–4k iterations (batch size 4, sequence length 4096; Table 5). The baseline at 1k steps (140% / 202% WER) frequently hallucinates transcripts and produces unreliable stopping behavior, consistent with observations in (Nguyen et al., 2025). Although performance improves with iterations, it remains far worse even at 4k (20% / 40% WER). In contrast, LST achieves 6.8% / 10.4% WER at 1k while reducing the context units during ASR inference. Under the same setup, we evaluate TTS reconstruction after 20k fine-tuning steps (Table 5). LST matches the baseline in CER and reduces the generation length by during TTS inference. CER is computed from Whisper-based transcriptions (Radford et al., 2023). Together, LST enables faster ASR adaptation and preserves TTS quality while requiring fewer context units and decoding steps at inference.
Visualization of Word-Level Speech Patch Embeddings